On July 2, 2019, Vicuna Inc. purchased equipment for $720,000. This equipment has an estimated useful life of six years and an estimated residual value of $30,000. Depreciation is taken for the portion of the year the asset is used. The asset is a Class 8 asset with a maximum CCA rate of 20%. Vicuna has a December year end.

Instructions

a) Complete the schedule below by determining the depreciation expense/CCA and year-end book values/UCC for 2019 and 2020 using the

1.double-declining-balance method.

2.capital cost allowance method (using maximum rate).

Double-Declining-Balance Method20192020

Depreciation expense for year

Accumulated depreciation

Year-end book value

Capital Cost Allowance Method 2019   2020

CCA for year

End of year UCC

Total CCA claimed

b) Instead, assume Vicuna used straight-line depreciation during 2019 and 2020. During 2021, the company determined that the equipment would be useful to the company for only one more year beyond 2021. Residual value is estimated at $40,000. Calculate the amount of depreciation expense for the 2021 income statement.
Answer #1
July 02,2019 Purchase value of Equipment= $                720,000
Salvage value $                  30,000
Useful Life in years =                                6
So SL depreciation rate per year =1/6= 16.67%
Double declining depreciation rate/year=2*Sl deprn= 33.33%
As the Equipment is purchased in July
we are using half year convention for all
the methods.
Ans a.
1. Depreciation by DD method
Year DD Depreciation for the year Accumulated Depreciation Year End Book Value
2019 $          120,000.00 $          120,000.00 $                                    600,000
2020 $          200,000.00 $          320,000.00 $                                    400,000
2. Depreciation by CCA 20%
Year CCA for the year End of Year UCC Total CCA aclaimed
2019 $             72,000.00 $          648,000.00 $                                      72,000
2020 $          129,600.00 $          518,400.00 $                                    201,600
In both the cases half year convention followed for the first year and depreciation rate applied on the reducing
asset book balance of previous year end.
Ans b.
July 02,2019 Purchase value of Equipment= $                720,000
Salvage value $                  30,000
Depreciable value $                690,000
Useful Life in years =                                6
So SL depreciation rate per year =1/6= 16.67%
Annual Depreciation =690000/6=                    115,000
Depreciation expense for 2019 (half year convention)=                      57,500
Depreciation expense for 2020=                    115,000
Total Depreciation cliamed in 2019+2020=                    172,500
Asset Book value on Jan 1,2021=720,000-172500=                    547,500
Revised Salvage estimate                      40,000
Revised depreciable value =                    507,500
Revised useful life in years =                                2
Revised Annual depreciation =                    253,750
So Depreciation expense for 2021 Income statement= $          253,750.00